• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Saha

    Log entry time 07:58:26 on January28,2002

    Entry number 81076

    keyword=Shift Summary

    Shift Summary
    ______________

    Midnight: Run 1798, KIN 4E, Run E*, LH2+Radiator (Cont.)
    00:43 Run 1798 ended, 300k events
    Tested Arc scanners. Scan #1483
    ep (beamline) LeCroy HV tripped. Restarted them.
    00:51 Run 1799 started, KIN 4E, Run 4E* (cont.)
    01:50 Coda Crash. Run 1799 finished, 130 k events
    02:23 Run 1800 started. Optics target run. 30 uA, no raster
    02:37 Run 1800 ended. 114 k events
    02:40 Starting arc-energy measurement
    04:10 Finished arc energy measurement
    Arc Energy = 4615.7 MeV +/- 0.9 MeV (HALOG 81074)
    Tiefenback Energy = 4615.5 MeV +/- 4.5 MeV
    Trouble moving left bogie. Rebooted Bogie IOC. Still did not work.
    05:10 Ask for access to reset power supplies on bogies. This should work according to Mark Stevens (Teck on call). Didn't work. Mark came in and is working on it at end of shift